OpenStudy (anonymous):

how do you evaluate this fractional expression: (3/4)^3

OpenStudy (saifoo.khan):

\[\frac{3^3}{4^3}=\frac{27}{64}\]

OpenStudy (anonymous):

thank you (:

OpenStudy (saifoo.khan):

you welcome! [: